5. Convert LS** Idzdydr into an integral in spherical coordinates.
Answer #1

5.) 1/4 x2 54x2y2 1 dzdxdx T = 1-2 limit of Integral 0<Z74-32-42 OsYE √4-x2 -2.252 Lef. z=plast =poso. Sind, y=p Soro. Sind ,